OpenStudy (anonymous):

1. How are the functions y = x and y = x – 3 related? How are their graphs related? Each output for y = x – 3 is 3 more than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated down 3 units. Each output for y = x – 3 is 3 more than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated up 3 units. Each output for y = x – 3 is 3 less than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated down 3 units. Each output for y = x – 3 is 3 less than the corresponding output f

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@mathstudent55

zepdrix (zepdrix):

Just test a number. For x=4:\[\large\rm y=x-3\qquad\to\qquad y=1\]\[\large\rm y=x\qquad\qquad\to\qquad y=4\] So what can you say about the outputs? :o

OpenStudy (anonymous):

they are different

OpenStudy (anonymous):

im confused tho.. what does that have to do with the problem? @zepdrix

zepdrix (zepdrix):

Each output of y=x-3 is `3 more` or `3 less` than y=x. We're picking an x, and comparing the outputs. This will help us narrow down our options. When we plugged in x=4, Did we get y=x-3 output being `3 larger` or `3 smaller` than the y=x output

zepdrix (zepdrix):

You can't just say that they're different, that's not enough information :)

OpenStudy (anonymous):

ohhhhh. that makes sense. thanks for your help

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

@confluxepic